OpenStudy (please.help.me):

tan2 x + tan x − 12 = 0 Use inverse functions where needed to find all solutions of the equation in the interval [0, 2π). (Enter your answers as a comma-separated list.)

OpenStudy (please.help.me):

Screenshot of the math problem>

OpenStudy (steve816):

Heyyy, looks like the above equation can be easily factored! Can you try factoring it?

OpenStudy (please.help.me):

Screenshot of example of what the answer is supposed to look like

OpenStudy (please.help.me):

I am confused on how to factor it because I know 1 x -30 = -30(The last term). But the number that are supposed to be added together have to equal 1! tan2 x + tan x − 12 = 0 (tan -1 )(tan +30 )

OpenStudy (steve816):

Not quite right! This would be the factored form\[\huge(\tan(x)+4)(\tan(x)-3)\]

OpenStudy (please.help.me):

I'm sorry! I accidentally copied the example from the screenshot on my paper!

OpenStudy (please.help.me):

I was using tan2 x + tan x − 30 = 0 lol

OpenStudy (please.help.me):

How would you factor that though?^

OpenStudy (steve816):

Use +6 and -5

OpenStudy (steve816):

\[\large (\tan x+6)(\tan x -5)\]

OpenStudy (sshayer):

\[\tan ^2x+\tan x-12=0\] \[\tan ^2 x+4\tan x-3\tan x-12=0\] \[\tan x(\tan x+4)-3(\tan x+4)=0\] \[\left( \tan x+4 \right)\left( \tan x-3 \right)=0\] \[tanx=-4\] Hence x lies in 2nd and 4th quadrant. \[x=\pi- \tan^{-1} (4),2 \pi-\tan^{-1}( 4)\] \[\tan x=3, x~ lies~ \in~ 1st~ and~ 3rd~quadrant.\] \[x=\tan^{-1} (3),\pi+\tan^{-1} (3)\]

OpenStudy (steve816):

^ nice

OpenStudy (please.help.me):

Thank you! Sorry I disappeared for a second!

OpenStudy (sshayer):

yw